OpenStudy (anonymous):

Prove by Induction: \[\sum_{i=1}^{n} i/2^{i} =2-((n+2)/2^{n})\] Prove by Induction: \[\sum_{i=1}^{n} i/2^{i} =2-((n+2)/2^{n})\] @Mathematics

OpenStudy (anonymous):

I just can't seem to find the inductive conclusion

OpenStudy (anonymous):

what you need to show is that \[2-\frac{n+2}{2^n}+\frac{n+1}{2^{n+1}}=2-\frac{n+3}{2^{n+1}}\] right?i think

OpenStudy (anonymous):

is that what you have?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so i guess it is algebra from here on in. i didn't do it, i just wrote it. let me see what i get

OpenStudy (anonymous):

I just can't get it to simplify and equal out

OpenStudy (anonymous):

yeah i get a numerator of \[3n+5\]!!

OpenStudy (anonymous):

that's what I get too

OpenStudy (anonymous):

but if it was minus instead of plus it would give the right thing

OpenStudy (anonymous):

hold on i think maybe this formula is not quite right

OpenStudy (anonymous):

i am a moron! totally forgot arithmetic!

OpenStudy (anonymous):

what is our error?

OpenStudy (anonymous):

our error is a really annoying one. we both thought that 2 - 3 + 5 = 2 - 8 !

OpenStudy (anonymous):

oh man

OpenStudy (anonymous):

in other words we forgot that we are subtracting, not adding!

OpenStudy (anonymous):

they are going to take away my math license now for sure. i bet it is clear now yes? you need \[\frac{n+1}{2^{n+1}}-\frac{n+2}{2^n}\] and i bet it works out right.

OpenStudy (kirbykirby):

yes hehe don't forget to carry out the minus sign!!

OpenStudy (anonymous):

wait hold on, what about the 2?

OpenStudy (anonymous):

it works fine. you want \[2-\frac{n+2}{2^n}+\frac{n+1}{2^{n+1}}=2-\frac{n+3}{2^{n+1}}\] the 2 in front just stays there.

OpenStudy (anonymous):

ahhh I see lol

OpenStudy (anonymous):

just make sure to compute the second part correctly. it is \[\frac{n+1}{2^{n+1}}-\frac{n+2}{2^n}\] \[\frac{n+1-2(n-2)}{2^{n+1}}\] \[\frac{-n-3}{2^{n+1}}\] \[-\frac{n+3}{2^{n+1}}\] just as you want it

OpenStudy (anonymous):

typo there , line 2 should be \[\frac{n+1-2(n+2)}{2^{n+1}}\]

OpenStudy (anonymous):

yeah, I've got it thanks This problem just tricks you into forgetting simple arithmetic Couldn't figure out what was going wrong for a while lol

OpenStudy (anonymous):

the old "order of operations" strikes again. please excuse my senile old aunt